### v4

# The distribution on permutations induced by a random parking function

Abstract.
A parking function on $[n]$ creates a permutation in $S_{n}$ via the order in which the $n$ cars appear in the $n$ parking spaces. Placing the uniform probability measure on the set of parking functions on $[n]$ induces a probability measure on $S_{n}$. We initiate a study of some properties of this distribution. Let $P_{n}^{\text{park}}$ denote this distribution on $S_{n}$ and let $P_{n}$ denote the uniform distribution on $S_{n}$. In particular, we obtain an explicit formula for $P_{n}^{\text{park}}(\sigma)$ for all $\sigma\in S_{n}$. Then we show that for all but an asymptotically $P_{n}$-negligible set of permutations, one has $P_{n}^{\text{park}}(\sigma)\in\left(\frac{(2-\epsilon)^{n}}{(n+1)^{n-1}},\frac{(2+\epsilon)^{n}}{(n+1)^{n-1}}\right)$. However, this accounts for only an exponentially small part of the $P_{n}^{\text{park}}$-probability. We also obtain an explicit formula for $P_{n}^{\text{park}}(\sigma^{-1}_{n-j+1}=i_{1},\sigma^{-1}_{n-j+2}=i_{2},\cdots,\sigma^{-1}_{n}=i_{j})$, the probability that the last $j$ cars park in positions $i_{1},\cdots,i_{j}$ respectively, and show that the $j$-dimensional random vector $(n+1-\sigma^{-1}_{n-j+l},n+1-\sigma^{-1}_{n-j+2},\cdots,n+1-\sigma^{-1}_{n})$ under $P_{n}^{\text{park}}$ converges in distribution to a random vector $(\sum_{r=1}^{j}X_{r},\sum_{r=2}^{j}X_{r},\cdots,X_{j-1}+X_{j},X_{j})$, where $\{X_{r}\}_{r=1}^{j}$ are IID with the Borel distribution. We then show that in fact for $j_{n}=o(n^{\frac{1}{6}})$, the final $j_{n}$ cars will park in increasing order with probability approaching 1 as $n\to\infty$.
Key words and phrases:
parking function, random permutation, Borel distribution
2010 Mathematics Subject Classification:
60C05, 05A05
1. Introduction and Statement of Results
Consider a row of $n$ parking spaces on a one-way street. A line of $n$ cars, numbered from 1 to $n$, attempt to park, one at a time. The $i$th car’s preferred space is spot number $\pi_{i}\in[n]$. If this space is already taken, then car $i$ proceeds forward and parks in the first available space, if one exists. If the car is unable to park, it exits the street. A sequence $\pi=\{\pi_{i}\}_{i=1}^{n}$ is called a parking function on $[n]$ if all $n$ cars are able to park. It is easy to see that $\pi$ is a parking function if and only if |$|\{i:\pi_{i}\leq j\}|\geq j$, for all $j\in[n]$. Let $\mathcal{P}_{n}$ denote the set of parking functions. It is well-known that $|\mathcal{P}_{n}|=(n+1)^{n-1}$. There are a number of proofs of this result; a particularly elegant one due to Pollack can be found in [3]. There is a large literature on parking functions and their generalizations; see, for example, the survey [5].
We can consider a random parking function by placing the uniform probability measure on $\mathcal{P}_{n}$. Denote this probability measure by $P^{\mathcal{P}_{n}}$. A study of random parking functions was initiated by Diaconis and Hicks in [1]. Since each parking function yields a permutation $\sigma=\sigma_{1}\cdots\sigma_{n}\in S_{n}$, where $\sigma_{j}$ is the number of the car that parked in space $j$, it follows that a random parking function induces a distribution on the set $S_{n}$ of permutations of $[n]$. In this paper we initiate a study of this distribution.
We will use the notation $P_{n}$ and $E_{n}$ to denote the uniform probability measure and the corresponding expectation on $S_{n}$. We will denote by $P_{n}^{\text{park}}$ the probability measure on $S_{n}$ induced by a random parking function in $\mathcal{P}_{n}$. The corresponding expectation will be denoted by $E_{n}^{\text{park}}$. To be more precise concerning the definition of the induced probability measure, define $T_{n}:\mathcal{P}_{n}\to S_{n}$ by $T_{n}(\pi)=\sigma$, if when using the parking function $\pi$, $\sigma_{j}$ is the number of the car that parked in space $j$, for $j\in[n]$. For example, if $n=4$ and $\pi=2213\in\mathcal{P}_{4}$, then we have $T_{4}(2213)=3124\in S_{4}$. We define
(1.1) $P_{n}^{\text{park}}(\sigma)=P^{\mathcal{P}_{n}}\left(T_{n}^{-1}(\{\sigma\})\right).$
For $1\leq i\leq n<\infty$ and $\sigma\in S_{n}$, define
(1.2) $l_{n,i}(\sigma)=\max\{l\in[i]:\sigma_{i}=\max(\sigma_{i},\sigma_{i-1},\cdots,\sigma_{i-l+1}\}.$
For $n\in\mathbb{N}$, define
$L_{n}(\sigma)=\prod_{i=1}^{n}l_{n,i}(\sigma),\ \sigma\in S_{n}.$
For example, if $\sigma=379218645\in S_{9}$, then $l_{n,i}(\sigma)=1$, for $i\in\{1,4,5,7,8\}$, $l_{n,i}(\sigma)=2$, for $i\in\{2,9\}$ and $l_{n,i}(\sigma)=3$, for $i\in\{3,6\}$. Thus, $L_{9}(\sigma)=1^{5}2^{2}3^{2}=36$. At certain points in the paper, it will be convenient to use the following variant of $l_{n,i}$. Define
(1.3) $\tilde{l}_{n,i}(\sigma)=l_{n,\sigma^{-1}_{i}}(\sigma).$
For example, if $\sigma=379218645$, then $\tilde{l}_{n,i}(\sigma)=1$, for $i\in\{1,2,3,4,6\}$, $\tilde{l}_{n,i}=2$, for $i\in\{5,7\}$ and $\tilde{l}_{n,i}=3$, for $i\in\{8,9\}$. Of course,
(1.4) $L_{n}(\sigma)=\prod_{i=1}^{n}l_{n,i}(\sigma)=\prod_{i=1}^{n}\tilde{l}_{n,i}(\sigma),\ \sigma\in S_{n}.$
We have the following theorem.
Theorem 1.
(1.5) $P_{n}^{\text{\rm park}}(\sigma)=\frac{L_{n}(\sigma)}{(n+1)^{n-1}},\ \sigma\in S_{n}.$
The following corollary of Theorem 1 is immediate, where the asymptotic behavior follows from Stirling’s formula, $n!\sim n^{n}e^{-n}\sqrt{2\pi n}$.
Corollary 1.
The expected value of the random variable $L_{n}=L_{n}(\sigma)$ on $(S_{n},P_{n})$ satisfies
(1.6) $E_{n}L_{n}=\frac{1}{n!}\sum_{\sigma\in S_{n}}L_{n}(\sigma)=\frac{(n+1)^{n-1}}{n!}\sim\frac{e^{n+1}}{\sqrt{2\pi}\thinspace n^{\frac{3}{2}}}.$
Although $E_{n}L_{n}$ is around $e^{n}$, with high $P_{n}$-probability $L_{n}(\sigma)$ is near $2^{n}$. We will prove the following weak convergence result for $L_{n}$.
Theorem 2.
For any $\epsilon\in(0,2]$, the random variable $L_{n}=L_{n}(\sigma)$ on $(S_{n},P_{n})$ satisfies
(1.7) $\lim_{n\to\infty}P_{n}((2-\epsilon)^{n}\leq L_{n}\leq(2+\epsilon)^{n}))=1.$
The following corollary follows immediately from Theorems 1 and 2.
Corollary 2.
For any $\epsilon\in(0,2]$, the random variable $P_{n}^{\text{\rm park}}=P_{n}^{\text{\rm park}}(\sigma)$ on $(S_{n},P_{n})$ satisfies
(1.8) $\lim_{n\to\infty}P_{n}\left(\bigg{\{}\sigma\in S_{n}:P_{n}^{\text{\rm park}}(\sigma)\in\left(\frac{(2-\epsilon)^{n}}{(n+1)^{n-1}},\frac{(2+\epsilon)^{n}}{(n+1)^{n-1}}\right)\bigg{\}}\right)=1.$
And the following corollary follows immediately from Theorem 1, Corollary 1 and Stirling’s formula.
Corollary 3.
The expectation of the random variable $P_{n}^{\text{\rm park}}=P_{n}^{\text{\rm park}}(\sigma)$ on $(S_{n},P_{n})$ is given by
(1.9) $E_{n}P_{n}^{\text{\rm park}}=\frac{1}{n!}\sim\frac{e^{n}}{\sqrt{2\pi}n^{n+\frac{1}{2}}}.$
Comparing (1.8) and (1.9), we see that for all but an asymptotically $P_{n}$-negligible set of permutations in $S_{n}$, the $P_{n}^{\text{park}}$-probability of a permutation in $S_{n}$ is approximately $\frac{2^{n}}{(n+1)^{n-1}}$, but the “average” $P_{n}^{\text{park}}$-probability of a permutation in $S_{n}$ is exponentially larger, namely asymptotic to $\frac{e^{n}}{\sqrt{2\pi}n^{n+\frac{1}{2}}}$. By Stirling’s formula, it also follows that there is a set $A_{n}\subset S_{n}$ with $P_{n}(A_{n})\to 1$ for which $P_{n}^{\text{park}}(A_{n})$ is around $(\frac{2}{e})^{n}$. There is an asymptotically $P_{n}$-negligible set of permutations in $S_{n}$ each of whose elements has super-exponentially larger $P_{n}^{\text{park}}$-probability than the average probability, and an asymptotically $P_{n}$-negligible set of permutations in $S_{n}$ each of whose elements has exponentially smaller $P_{n}^{\text{park}}$-probability than the average probability. In particular, we have the following corollary.
Corollary 4.
The maximum value of $P_{n}^{\text{\rm park}}=P_{n}^{\text{\rm park}}(\sigma)$ is equal to $\frac{n!}{(n+1)^{n-1}}\sim\frac{\sqrt{2\pi}\thinspace n^{\frac{3}{2}}}{e^{n+1}}$ and is attained uniquely at $\sigma=1\cdots n$. The minimum value of $P_{n}^{\text{\rm park}}$ is equal to $\frac{1}{(n+1)^{n-1}}$ and is attained uniquely at $\sigma=n\cdots 1$.
Proof.
The function $L_{n}=L_{n}(\sigma),\sigma\in S_{n}$, attains its maximum value $n!$ uniquely at $\sigma=1\cdots n$ and attains its minimum value 1 uniquely at $\sigma=n\cdots 1$. . ∎
Let $\sigma^{-1}$ denote the inverse permutation of $\sigma$. So $\sigma^{-1}_{k}=j$ if and only if $\sigma_{j}=k$. In car parking language,$\sigma^{-1}_{k}=j$ means that car number $k$ parked in space number $j$. From the definition of a parking function, it is obvious that
$P_{n}^{\text{park}}(\sigma^{-1}_{1}=j)=P^{\mathcal{P}_{n}}(\pi_{1}=j),\ j\in[n].$
In [1], the following asymptotic behavior was proven for $\pi_{1}$ (or any $\pi_{k}$ by symmetry):
(1.10) $\text{For fixed}\ j,P^{\mathcal{P}_{n}}(\pi_{1}=j)\sim\frac{1+P(X\geq j)}{n};$
$\text{For fixed}\ j,P^{\mathcal{P}_{n}}(\pi_{1}=n-j)\sim\frac{P(X\leq j+1)}{n},$
$\text{where}\ X\ \text{is a random variable satisfying}\ P(X=j)=e^{-j}\frac{j^{j-1}}{j!},\ j=1,2,\cdots.$
The distribution $\{e^{-j}\frac{j^{j-1}}{j!}\}_{j=1}^{\infty}$ is called the Borel distribution. It is not obvious that it is a distribution, that is, that it sums to 1. For more on this, see [1, p.135]. It follows that (1.10) also holds with $P^{\mathcal{P}_{n}}$ and $\pi_{1}$ replaced respectively by $P_{n}^{\text{park}}$ and $\sigma^{-1}_{1}$. It would be nice to obtain asymptotic results for $P_{n}^{\text{park}}(\sigma^{-1}_{j_{n}}=k_{n})$, for general $j_{n},k_{n}$. We pursue this direction when $j_{n}$ is near $n$, that is for the last cars to park.
Theorem 3.
(1.11) $P_{n}^{\text{park}}(\sigma^{-1}_{n-j+1}=i_{1},\sigma^{-1}_{n-j+2}=i_{2},\cdots,\sigma^{-1}_{n}=i_{j})=$
$\frac{(n-j)!\left(\prod_{l=1}^{j}(k_{l}-k_{l-1})^{k_{l}-k_{l-1}-2}\right)(n-k_{j}+1)^{n-k_{j}-1}\prod_{l=1}^{j}\tilde{l}_{n,n-j+l}(\sigma)}{(n+1)^{n-1}\left(\prod_{l=1}^{j}(k_{l}-k_{l-1}-1)!\right)(n-k_{j})!},$
where $\{k_{l}\}_{l=1}^{j}$ is the increasing rearrangement of $\{i_{l}\}_{l=1}^{j}$, $k_{0}=0$ and $\tilde{l}_{n,\cdot}(\sigma)$ is as in (1.3).
One can use Theorem 3 to make explicit calculations. Here we treat the cases $j=1$ and $j=2$ in depth in Corollaries 5-8. This will lead us to a result for general $j$ in Corollary 9 and for growing $j=j_{n}$ in Theorem 4.
We begin with the case $j=1$.
Corollary 5.
(1.12) $P_{n}^{\text{park}}(\sigma^{-1}_{n}=k)=\frac{1}{n}k^{k}\binom{n}{k}\frac{(n-k+1)^{n-k-1}}{(n+1)^{n-1}},\ k\in[n].$
Proof.
Immediate from (1.11), noting that if $\sigma^{-1}_{n}=k$, then $\tilde{l}_{n,n}(\sigma)=k$. ∎
From Corollary 5 we obtain the following asymptotic formulas.
Corollary 6.
i. For fixed $m\in\mathbb{N}$,
(1.13) $\lim_{n\to\infty}P_{n}^{\text{park}}(\sigma^{-1}_{n}=n+1-m)=\frac{m^{m-1}e^{-m}}{m!}.$
Thus, the random variable $n+1-\sigma^{-1}_{n}$ under $P_{n}^{\text{park}}$ converges in distribution to a random variable $X$ with the Borel distribution.
ii. For fixed $k\in\mathbb{N}$,
(1.14) $P_{n}^{\text{park}}(\sigma^{-1}_{n}=k)\sim\frac{k^{k}e^{-k}}{k!}\frac{1}{n},\ \text{as}\ n\to\infty.$
iii. Let $c_{n}n$ be an integer with $\lim_{n\to\infty}c_{n}=c\in(0,1)$. Then
(1.15) $P_{n}^{\text{park}}(\sigma^{-1}_{n}=c_{n}n)\sim\frac{1}{(2\pi c)^{\frac{1}{2}}(1-c)^{\frac{3}{2}}}\frac{1}{n^{\frac{3}{2}}}.$
Proof.
The proof follows from (1.12) and standard asymptotic analysis. We write out the proof of part (iii). Substituting in (1.12), we have
(1.16) $P_{n}^{\text{park}}(\sigma^{-1}_{n}=c_{n}n)=\frac{1}{n}(c_{n}n)^{c_{n}n}\frac{n!}{(c_{n}n)!((1-c_{n})n)!}\frac{(n-c_{n}n+1)^{n-c_{n}n-1}}{(n+1)^{n-1}}.$
Replacing the three factorials on the right hand side of (1.16) by their Stirling’s formula approximations, $m!\sim m^{m}e^{-m}\sqrt{2\pi m}$ as $m\to\infty$, and performing many cancelations, we have
$\frac{1}{n}(c_{n}n)^{c_{n}n}\frac{n!}{(c_{n}n)!((1-c_{n})n)!}\frac{(n-c_{n}n+1)^{n-c_{n}n-1}}{(n+1)^{n-1}}\sim$
$\frac{n^{n-1}\left(1+\frac{1}{n-c_{n}n}\right)^{n-c_{n}n-1}}{n\sqrt{c_{n}}(1-c_{n})\sqrt{2\pi(1-c_{n})n}\thinspace(n+1)^{n-1}}.$
Since $\lim_{n\to\infty}\frac{n^{n-1}\left(1+\frac{1}{n-c_{n}n}\right)^{n-c_{n}n-1}}{(n+1)^{n-1}}=1$, we conclude that the right hand side of (1.16) is asymptotic to $\frac{1}{(2\pi c)^{\frac{1}{2}}(1-c)^{\frac{3}{2}}}\frac{1}{n^{\frac{3}{2}}}$. ∎
Remark. One can check that $\frac{k^{k}e^{-k}}{k!}$ is decreasing in $k$. Thus, from part (ii), for fixed $k$, $P_{n}^{\text{park}}(\sigma^{-1}_{n}=k)$ is on the order $\frac{1}{n}$ and decreasing in $k$. From part (iii), the probability of $\sigma^{-1}_{n}$ taking on any particular value in the bulk is even smaller, namely on the order $n^{-\frac{3}{2}}$. And from part (i), the distance of $\sigma^{-1}_{n}$ from $n$ converges in distribution as $n\to\infty$.
We now turn to the case $j=2$.
Corollary 7.
Let
(1.17) $A_{n}(a,b)=\frac{(n-2)!a^{a-2}(b-a)^{b-a-2}(n-b+1)^{n-b-1}}{(a-1)!(b-a-1)!(n-b)!(n+1)^{n-1}},\ \text{for}\ 1\leq a<b\leq n.$
Then
(1.18) $P_{n}^{\text{park}}(\sigma^{-1}_{n-1}=l,\sigma^{-1}_{n}=m)=\begin{cases}lmA_{n}(l,m),\ \text{if}\ 1\leq l<m\leq n;\\ (l-m)mA_{n}(m,l),\ \text{if}\ 1\leq m<l\leq n.\end{cases}.$
Proof.
Immediate from (1.11), noting that if $\sigma^{-1}_{n-1}=l$ and $\sigma^{-1}_{n}=m$, then
$\begin{cases}\tilde{l}_{n,n-1}(\sigma)=l,\ \tilde{l}_{n,n}(\sigma)=m,\ \text{if}\ l<m;\\ \tilde{l}_{n,n-1}(\sigma)=l-m,\ \tilde{l}_{n,n}(\sigma)=m,\ \text{if}\ l>m.\end{cases}$
From Corollary 7 we obtain the following asymptotic formulas.
Corollary 8.
i. For fixed $l,m$,
(1.19) $\lim_{n\to\infty}P_{n}^{\text{park}}(\sigma^{-1}_{n-1}=n+1-l,\sigma^{-1}_{n}=n+1-m)=\frac{(l-m)^{l-m-1}m^{m-1}}{(l-m)!m!}e^{-l},\ 1\leq m<l;$
(1.20) $P_{n}^{\text{park}}(\sigma^{-1}_{n-1}=n+1-l,\sigma^{-1}_{n}=n+1-m)\sim\frac{(m-l)^{m-l}l^{l-1}}{l!(m-l)!}\frac{1}{n},\ 1\leq l<m.$
Thus, the random vector $(n+1-\sigma^{-1}_{n-1},n+1-\sigma^{-1}_{n})$ under $P_{n}^{\text{park}}$ converges in distribution to a random vector $(X_{1}+X_{2},X_{2})$ where $X_{1}$ and $X_{2}$ are IID with the Borel distribution.
ii. For fixed, $l,m$,
(1.21) $P_{n}^{\text{park}}(\sigma^{-1}_{n-1}=l,\sigma^{-1}_{n}=m)\sim\begin{cases}\frac{l^{l}}{l!}\frac{(m-l)^{m-l-1}m}{(m-l)!}e^{-m}\frac{1}{n^{2}},\ 1\leq l<m;\\ \frac{m^{m}}{m!}\frac{(l-m)^{l-m}}{(l-m)!}e^{-l}\frac{1}{n^{2}},\ 1\leq m<l.\end{cases}$
Proof.
Standard asymptotic analysis. ∎
Part (i) of Corollaries 6 and 8 lead us to the following result for all $j$, which we will prove.
Corollary 9.
Let $j\in\mathbb{N}$. The $j$-dimensional random vector
$(n+1-\sigma^{-1}_{n-j+l},n+1-\sigma^{-1}_{n-j+2},\cdots,n+1-\sigma^{-1}_{n})$ under $P_{n}^{\text{park}}$ converges in distribution to a random vector $(\sum_{r=1}^{j}X_{r},\sum_{r=2}^{j}X_{r},\cdots,X_{j-1}+X_{j},X_{j})$, where $\{X_{r}\}_{r=1}^{j}$ are IID with the Borel distribution.
In particular then, for any $j\in\mathbb{N}$, the last $j$ cars from among the $n$ cars will park in increasing order with probability approaching 1 as $n\to\infty$:
(1.22) $\lim_{n\to\infty}P_{n}^{\text{park}}(\sigma^{-1}_{n+1-j}<\sigma^{-1}_{n+1-j+2}<\cdots<\sigma^{-1}_{n})=1.$
In fact, we can extend (1.22) to the last $j_{n}$ cars, where $j_{n}=o(n^{\frac{1}{6}})$.
Theorem 4.
For $j_{n}=o(n^{\frac{1}{6}})$, the last $j_{n}$ cars from among the $n$ cars will park in increasing order with probability approaching 1 as $n\to\infty$:
(1.23) $\lim_{n\to\infty}P_{n}^{\text{park}}(\sigma^{-1}_{n+1-j_{n}}<\sigma^{-1}_{n+1-j_{n}+1}<\cdots<\sigma^{-1}_{n})=1.$
We will also prove the following result, without reference to Theorem 3.
Proposition 1.
For any $m\in[n]$,
(1.24) $P_{n}^{\text{park}}\left(\{\sigma_{1},\cdots,\sigma_{m}\}=[m]\right)=\left(\frac{m+1}{n+1}\right)^{m}.$
We have the following immediately corollary of Proposition 1.
Corollary 10.
For any $m\in\mathbb{N}$,
(1.25) $\lim_{n\to\infty}P_{n}^{\text{park}}(\{\sigma_{1},\cdots,\sigma_{n-m}\}=[n-m])=e^{-m}$
Remark. Corollary 10 can also be deduced from Corollary 9.
The proof of Theorem 1 is given in section 2 and the proof of Theorem 2 is given in section 3. The proof of Theorem 3 is given in section 4, the proof of Corollary 9 is given in section 5 and the proof of Theorem 4 is given in section 6. The proof of Proposition 1 is given in section 7.
2. Proof of Theorem 1
Recall the definitions of $l_{n,i}$ and $\tilde{l}_{n,i}$ in (1.2) and (1.3). For the proof of the theorem, it will be convenient to work with $\tilde{l}_{n,i}$. Recall from (1.4) that
$L_{n}(\sigma)=\prod_{i=1}^{n}\tilde{l}_{n,i}(\sigma).$
The theorem will follow if we show that for each $\sigma\in S_{n}$, there are $L_{n}(\sigma)$ different parking functions $\pi\in\mathcal{P}_{n}$ such that $T_{n}(\pi)=\sigma$, where $T_{n}$ is as in the paragraph containing equation (1.1). Before giving a formal proof of the theorem, we illustrate the proof with a concrete example, from which the general result should be clear. Consider the permutation $\sigma=379218645\in S_{9}$. We look for those $\pi\in\mathcal{P}_{9}$ that satisfy $T_{9}(\pi)=\sigma$. From the definition of the parking process and from the definition of $T_{n}$, we need $\pi_{1}=5$ in order to have $\sigma_{5}=1$, $\pi_{2}=4$ in order to have $\sigma_{4}=2$, $\pi_{3}=1$ in order to have $\sigma_{1}=3$ and $\pi_{4}=8$ in order to have $\sigma_{8}=4$. In order to have $\sigma_{9}=5$, we can either have $\pi_{5}=9$, in which case car number 5 parks in its preferred space 9, or alternatively, $\pi_{5}=8$, in which case car number 5 attempts to park in its preferred space 8 but fails, and then moves on to space 9 and parks. Then we need $\pi_{6}=7$ in order to have $\sigma_{7}=6$. Then similar to the explanation regarding $\pi_{5}$, we need $\pi_{7}$ to be either 1 or 2 in order to have $\sigma_{2}=7$. In order to have $\sigma_{6}=8$, we can have either $\pi_{8}=6$, in which car number 8 parks directly in its preferred space 6, or alternatively $\pi_{8}=5$, in which case car number 8 tries and fails to park in space number 5 and then parks in space number 6, or alternatively, $\pi_{8}=4$, it which case car number 8 tries and fails to park in space number 4 and then also in space number 5, before finally parking in space number 6. Similarly, we need $\pi_{9}$ to be equal to 1,2 or 3 in order to have $\sigma_{9}=3$. Thus, there are $1\times 1\times 1\times 1\times 2\times 1\times 2\times 3\times 3=\prod_{i=1}^{9}\tilde{l}_{9,i}(\sigma)=L_{9}(\sigma)$ different parking functions $\pi\in\mathcal{P}_{9}$ that yield $T_{9}(\pi)=\sigma$.
To give a formal proof for the general case, fix $\sigma\in S_{n}$. In order to have $T_{n}(\pi)=\sigma$, first we need $\pi_{1}=\sigma^{-1}_{1}$. Thus there is just one choice for $\pi_{1}$, and note that $\tilde{l}_{n,1}(\sigma)=1$. Now let $k\in[n-1]$ and assume that we have chosen $\pi_{1},\cdots\pi_{k}$ in such a way that car number $i$ has parked in space $\sigma^{-1}_{i}$, for $i\in[k]$. We now want car number $k+1$ to park in space $\sigma^{-1}_{k+1}$. By construction, this space is vacant at this point, and so are the $\tilde{l}_{n,k+1}(\sigma)-1$ spaces immediately to the left of this space. However the space $\tilde{l}_{n,k+1}$ spaces to the left of this space is not vacant (or possibly this space doesn’t exist–it would be the zeroth space). Thus, by the parking process, car number $k+1$ will park in space $\sigma^{-1}_{k+1}$ if and only if $\pi_{k+1}$ is equal to one of the $\tilde{l}_{n,k+1}(\sigma)$ numbers $\sigma^{-1}_{k+1},\sigma^{-1}_{k+1}-1,\cdots,\sigma^{-1}_{k+1}-\tilde{l}_{n,k+1}(\sigma)+1$. This shows that there are $L_{n}(\sigma)=\prod_{i=1}^{n}\tilde{l}_{n,i}(\sigma)$ different parking functions $\pi$ satisfying $T_{n}(\pi)=\sigma$. $\square$
3. Proof of Theorem 2
We begin with several preliminary results. Recall that $P_{n}$ is the uniform probability measure on $S_{n}$.
Lemma 1.
(3.1) $P_{n}(l_{n,i}=j)=\begin{cases}\frac{1}{j}-\frac{1}{j+1}=\frac{1}{j(j+1)},\ j=1,\cdots,i-1;\\ \frac{1}{i},\ j=i.\end{cases}.$
Proof.
Fix $i$ and let $j\in[i]$. The event $\{l_{n,i}(\sigma)\geq j\}$ is the event $\big{\{}\sigma_{i}=\max\{\sigma_{i},\sigma_{i-1},\cdots,\sigma_{i-j+1}\}\big{\}}$. Since $P_{n}$ is the uniform distribution on $S_{n}$, we have
[…]
The lemma now follows. ∎
We now write
(3.3) $\mathcal{S}_{n}:=\log L_{n}=\sum_{i=1}^{n}\log l_{n,i}.$
From Lemma 1, we have
(3.4) $E_{n}\log l_{n,i}=\sum_{j=1}^{i-1}\frac{\log j}{j(j+1)}+\frac{\log i}{i}.$
Note that $E_{n}\log l_{n,i}$ does not depend on $n$, but of course it is only defined for $1\leq i\leq n$.
Lemma 2.
(3.5) $\lim_{n,i\to\infty}E_{n}\log l_{n,i}=\log 2.$
Proof.
Recall the Abel-type summation formula [4]:
[…]
We apply this formula with $a(r)=\frac{1}{r(r+1)}=\frac{1}{r}-\frac{1}{r+1}$ and $f(r)=\log r$. We have $A(r)=1-\frac{1}{r+1}=\frac{r}{r+1}$. Recalling (3.4), we obtain
[…]
$\lim_{i\to\infty}\left(\frac{i}{i+1}\log i-\log(i+1)+\log 2\right)=\lim_{i\to\infty}\left(\log\frac{i}{i+1}-\frac{\log i}{i+1}+\log 2\right)=\log 2.$
From (3.3) and (3.5), we conclude that
(3.6) $\lim_{n\to\infty}\frac{E_{n}\mathcal{S}_{n}}{n}=\lim_{n\to\infty}\frac{1}{n}E_{n}\log L_{n}=\log 2.$
We now consider $E_{n}\mathcal{S}_{n}^{2}$. We have
(3.7) $E_{n}\mathcal{S}_{n}^{2}=E_{n}\left(\sum_{i=1}^{n}\log l_{n,i}\right)^{2}=\sum_{i=1}^{n}E_{n}\log l_{n,i}^{2}\thinspace+2\sum_{1\leq i<j\leq n}E_{n}\log l_{n,i}\log l_{n,j}.$
We have the following proposition.
Proposition 2.
For $1\leq i<j\leq n$, the random variables $l_{n,i}$ and $l_{n,j}$ on $(S_{n},P_{n})$ are negatively correlated; that is,
(3.8) $P_{n}(l_{n,i}\geq k,\l_{n,j}\geq k)\leq P_{n}(l_{n,i}\geq k)P_{n}(l_{n,j}\geq l),\ \text{for}\ k,l\geq 1.$
Proof.
Since $P_{n}$ is the uniform probability measure on $S_{n}$, for any $k\leq i$, the events $\{l_{n,i}\geq k\}=\big{\{}\sigma_{i}=\max(\sigma_{i},\cdots,\sigma_{i-k+1})\big{\}}$ and $\{l_{n,j}\geq l\}=\big{\{}\sigma_{j}=\max(\sigma_{j},\cdots,\sigma_{j-l+1})\big{\}}$ are independent if $l\leq j-i$. Thus, (3.8) holds with equality in these cases.
Consider now the case $k\leq i$ and $j-i+1\leq l\leq j$. In this case
(3.9) $\{l_{n,i}\geq k,l_{n,j}\geq l\}=\big{\{}\sigma_{j}=\max(\sigma_{j},\sigma_{j-1},\cdots,\sigma_{r})\big{\}}\cap\big{\{}\sigma_{i}=\max(\sigma_{i},\sigma_{i-1},\cdots,\sigma_{i-k+1})\big{\}},$
$\text{where}\ r=\min(i-k+1,j-l+1).$
We have
(3.10) $P_{n}\left(\sigma_{j}=\max(\sigma_{j},\sigma_{j-1},\cdots,\sigma_{r})\right)=\frac{1}{\max(l,j-i+k)}\leq\frac{1}{l}.$
Also,
(3.11) $P_{n}\left(\sigma_{i}=\max(\sigma_{i},\sigma_{i-1},\cdots,\sigma_{i-k+1}|\sigma_{j}=\max(\sigma_{j},\sigma_{j-1},\cdots,\sigma_{r})\right)=\frac{1}{k},$
because $\{\sigma_{i},\cdots,\sigma_{i-k+1}\}\subset\{\sigma_{j},\sigma_{j-1},\cdots,\sigma_{r}\}$. The proposition follows from (3.9)–(3.11) and (3.2). ∎
We can now prove the theorem.
Proof of Theorem 2. Since $l_{n,i}$ and $l_{n,j}$ are negatively correlated, one has $E_{n}f(l_{n,i})g(l_{n,j})\leq E_{n}f(l_{n,i})E_{n}g(l_{n,j})$, if $f$ and $g$ are increasing functions on $[n]$. In particular then,
(3.12) $E_{n}\log l_{n,i}\log l_{n,j}\leq E_{n}\log l_{n,i}E_{n}\log l_{n,j}.$
From (3.3) and (3.12), a standard straightforward calculation gives
(3.13) $\text{\rm Var}(\mathcal{S}_{n})\leq\sum_{i=1}^{n}\text{\rm Var}(\log l_{n,i}).$
From (3.1), we have
$E_{n}(\log l_{n,i})^{2}=\sum_{j=1}^{i-1}\frac{(\log j)^{2}}{j(j+1)}+\frac{(\log i)^{2}}{i(i+1)}.$
Using this with (3.5) and (3.13), we conclude that there exists a $C>0$ such that
(3.14) $\text{\rm Var}(\mathcal{S}_{n})\leq Cn,\ n\in\mathbb{N}.$
From (3.6) and(3.14), it follows from the second moment method (Chebyshev’s inequality) that
(3.15) $\lim_{n\to\infty}P_{n}(\log 2-\epsilon\leq\frac{\mathcal{S}_{n}}{n}\leq\log 2+\epsilon)=1,\ \text{for all}\ \epsilon>0.$
Now (1.7) follows from (3.15) and (3.3). $\square$
4. Proof of Theorem 3
We count how many parking functions $\{\pi_{l}\}_{l=1}^{n}$ yield the event $\{\sigma^{-1}_{n-j+1}=i_{1},\sigma^{-1}_{n-j+2}=i_{2},\cdots,\sigma^{-1}_{n}=i_{j}\}$. In order for this event to occur, the first $n-j$ cars have to park arbitrarily in the spots $[n]-\{k_{l}\}_{l=1}^{j}$. Then the last $j$ cars have to park, with each one in its appropriate space.
We now analyze the parking of the first $n-j$ cars. In order for $k_{1}-1$ cars from among the first $n-j$ cars to park in the first $k_{1}-1$ spaces, but for no one of the first $n-j$ cars to park in the $k_{1}$th space, it follows from the definition of the parking process that there must be a collection of $k_{1}-1$ of the $\{\pi_{i}\}_{l=1}^{n-j}$ which constitute a parking function of length $k_{1}-1$. We can freely choose which of the first $n-j$ cars to use for these spaces. Since there are $k_{1}^{k_{1}-2}$ parking functions of length $k_{1}-1$, this gives $\binom{n-j}{k_{1}-1}k_{1}^{k_{1}-2}$ different choices.
In order for $k_{2}-k_{1}-1$ cars, from among the $n-j-(k_{1}-1)$ cars that still remain from among the first $n-j$ cars, to park in the spaces $k_{1}+1,\cdots,k_{2}-1$, but for no one of the first $n-j$ cars to park in the $k_{1}$th space or the $k_{2}$th space, there must be a collection of $k_{2}-k_{1}-1$ of the remaining $n-j-(k_{1}-1)$ members of $\{\pi_{l}\}_{l=1}^{n-j}$ that constitute a parking function of length $k_{2}-k_{1}-1$, but shifted forward by $k_{1}$ spaces. (If $\{\pi_{l}\}_{l=1}^{m}$ is a parking function of length $m$, then we call the collection $\{a+\pi_{l}\}_{l=1}^{m}$, where $a\in\mathbb{N}$, a parking function of length $m$ shifted forward by $a$ spaces. Obviously, the number of such parking functions coincides with the number of parking functions of length $m$.) We can freely choose which of the remaining $n-j-(k_{1}-1)$ cars from among the first $n-j$ cars to use for these spaces. Thus, there are $\binom{n-j-(k_{1}-1)}{k_{2}-k_{1}-1}(k_{2}-k_{1})^{k_{2}-k_{1}-2}$ different choices.
We continue counting in this fashion until we have chosen cars for the $k_{j}-k_{j-1}-1$ spaces between spaces $k_{j-1}$ and $k_{j}$. Recall that $k_{0}$ has been defined to be 0. Now we need for all of the $n-j-\sum_{i=1}^{j}(k_{i}-k_{i-1}-1)=n-k_{j}$ cars that remain from among the first $n-j$ cars to park in the spaces $k_{j}+1,\cdots,n$. In order for this to occur, but for no one of the first $n-j$ cars to park in the $k_{j}$th space, these remaining members of $\{\pi_{i}\}_{l=1}^{n-j}$ must constitute a parking function of length $n-k_{j}$, but shifted forward by $k_{j}$ spaces. This gives $(n-k_{j}+1)^{n-k_{j}-1}$ different choices.
Thus, we have shown that the number of ways to choice $\{\pi_{l}\}_{l=1}^{n-j}$ is
(4.1) $\binom{n-j}{k_{1}-1}k_{1}^{k_{1}-2}\binom{n-j-k_{1}+1}{k_{2}-k_{1}-1}(k_{2}-k_{1})^{k_{2}-k_{1}-2}\cdots\times$
$\binom{n-j-\sum_{l=1}^{j-1}(k_{l}-k_{l-1})}{k_{j}-k_{j-1}-1}(k_{j}-k_{j-1})^{k_{j}-k_{j-1}-2}(n-k_{j}+1)^{n-k_{j}-1}=$
$\frac{(n-j)!}{(n-k_{j})!\prod_{l=1}^{j}(k_{l}-k_{l-1}-1)!}\thinspace(n-k_{j}+1)^{n-k_{j}-1}\prod_{l=1}^{j}(k_{l}-k_{l-1})^{k_{l}-k_{l-1}-2}.$
Now we count how many ways we can choice the $\{\pi_{l}\}_{l=n-j+1}^{n}$ in order that $\sigma_{n-j+l}=i_{l},\ l=1,\cdots,j$, or equivalently, in order that car $n-j+l$ park in space $i_{l},\ l=1,\cdots,j$. This calculation uses the same reasoning that was used in the proof of Theorem 1. After the first $n-j$ cars have parked appropriately as above, the spaces $\{k_{l}\}_{l=1}^{j}=\{i_{l}\}_{l=1}^{j}$ are still vacant. From the definition of the parking process and from the definition of $\tilde{l}_{n,n-j+1}(\sigma)$, in order that car $n-j+1$ park in space $i_{1}$, $\pi_{n-j+1}$ must take on one of the values $i_{1}.i_{1}-1,\cdots i_{1}-\tilde{l}_{n,n-j+1}(\sigma)+1$. Then in order that car $n-j+2$ park in space $i_{2}$, $\pi_{n-j+2}$ must take on one of the values $i_{2},i_{2}-1,\cdots i_{2}-\tilde{l}_{n,n-j+2}(\sigma)+1$. Continuing like this, we conclude that there are $\prod_{l=1}^{j}\tilde{l}_{n,n-j+l}(\sigma)$ choices. Using this with (4.1), and noting that there are $(n+1)^{n-1}$ different parking functions of length $n$, we obtain (1.11). $\square$
5. Proof of Corollary 9
Let $\{m_{r}\}_{r=1}^{j}\subset\mathbb{N}$. By Theorem 3, for $n\geq\sum_{r=1}^{j}m_{r}$, the probability
$P_{n}^{\text{park}}(n+1-\sigma^{-1}_{n-j+l}=\sum_{r=l}^{j}m_{r};\ l=1,\cdots,j)$ is given by the formula on the right hand side of (1.11) with with $i_{l}=n+1-\sum_{r=l}^{j}m_{r},\ l\in[j]$. Since these $\{i_{l}\}_{l=1}^{j}$ are increasing, we have $k_{l}=i_{l},\ l\in[j]$. So we have
(5.1) $k_{l}-k_{l-1}=m_{l-1},\ l=2,\cdots j;$
$k_{1}-k_{0}=n+1-\sum_{r=1}^{j}m_{r}.$
Since the $\{i_{l}\}_{l=1}^{j}$ are increasing and represent the positions respectively of the numbers $\{n-j+l\}_{l=1}^{j}$ in the permutation $\sigma$, it follows from the definition of $\tilde{l}_{n,\cdot}(\sigma)$ in (1.3) that
(5.2) $\tilde{l}_{n,n-j+l}(\sigma)=i_{l}=n+1-\sum_{r=l}^{j}m_{r}.$
Substituting from (5.1) and (5.2) into the right hand side of (1.11), the expression there becomes
(5.3) $\frac{(n-j)!\left(n+1-\sum_{r=1}^{j}m_{r}\right)^{n-1-\sum_{r=1}^{j}m_{r}}\left(\prod_{l=2}^{j}m_{l-1}^{m_{l-1}-2}\right)m_{j}^{m_{j}-2}\prod_{l=1}^{j}\left(n+1-\sum_{r=l}^{j}m_{r}\right)}{(n+1)^{n-1}(n-\sum_{r=1}^{j}m_{r})!\left(\prod_{l=2}^{j}(m_{l-1}-1)!\right)(m_{j}-1)!}.$
We rewrite the terms in (5.3) that do not involve $n$ as
(5.4) $\frac{\left(\prod_{l=2}^{j}m_{l-1}^{m_{l-1}-2}\right)m_{j}^{m_{j}-2}}{\left(\prod_{l=2}^{j}(m_{l-1}-1)!\right)(m_{j}-1)!}=\prod_{l=1}^{j}\frac{m_{l}^{m_{l}-1}}{m_{l}!}.$
The terms in (5.3) that involve $n$ satisfy
(5.5) $\lim_{n\to\infty}\frac{(n-j)!\left(n+1-\sum_{r=1}^{j}m_{r}\right)^{n-1-\sum_{r=1}^{j}m_{r}}\prod_{l=1}^{j}\left(n+1-\sum_{r=l}^{j}m_{r}\right)}{(n+1)^{n-1}(n-\sum_{r=1}^{j}m_{r})!}=$
$e^{-\sum_{r=1}^{j}m_{r}}.$
From (5.3)-(5.5), it follows that the right hand side of (1.11) with $i_{l}=n+1-\sum_{r=l}^{j}m_{r},\ l\in[j]$, converges as $n\to\infty$ to $\prod_{l=1}^{j}\frac{m_{l}^{m_{l}-1}e^{-m_{l}}}{m_{l}!}$. Thus, we have proven that
(5.6) $\lim_{n\to\infty}P_{n}^{\text{park}}(n+1-\sigma^{-1}_{n-j+l}=\sum_{r=l}^{j}m_{r};\ l=1,\cdots,j)=\prod_{l=1}^{j}\frac{m_{l}^{m_{l}-1}e^{-m_{l}}}{m_{l}!}.$
The corollary follows from this. $\square$
6. Proof of Theorem 4
From the proof of Corollary 9 up to (5.4), in particular, from (5.3) and (5.4), we have for $j_{n}\leq n$,
(6.1) $P_{n}^{\text{park}}\left(n+1-\sigma^{-1}_{n-j_{n}+1}=\sum_{r=1}^{j_{n}}m_{r},n+1-\sigma^{-1}_{n-j_{n}+2}=\sum_{r=2}^{j_{n}}m_{r},\cdots,n+1-\sigma^{-1}_{n}=m_{j_{n}}\right)=$
$\left(\prod_{r=1}^{j_{n}}\frac{m_{r}^{m_{r}-1}e^{-m_{r}}}{m_{r}!}\right)A(n,j_{n};m_{1}\cdots,m_{j_{n}})B(n,j_{n};m_{1},\cdots,m_{j_{n}}),\ \text{for}\ \sum_{r=1}^{j_{n}}m_{r}\leq n,\ m_{r}\geq 1,$
where
(6.2) $A(n,j;m_{1}\cdots,m_{j})=e^{\sum_{r=1}^{j}m_{r}}\left(\frac{n+1-\sum_{r=1}^{j}m_{r}}{n+1}\right)^{n-1}=e^{\sum_{r=1}^{j}m_{r}}\left(1-\frac{\sum_{r=1}^{j}m_{r}}{n+1}\right)^{n-1};$
$B(n,j;m_{1},\cdots,m_{j})=\frac{(n-j)!}{(n-\sum_{r=1}^{j}m_{r})!}\frac{\prod_{r=1}^{j}(n+1-\sum_{r=l}^{j}m_{r})}{(n+1-\sum_{r=1}^{j}m_{r})^{\sum_{r=1}^{j}m_{r}}}.$
By Stirling’s formula, $\frac{m^{m-1}e^{-m}}{m!}\sim\frac{1}{\sqrt{2\pi}}m^{-\frac{3}{2}}$. Thus, there exists a $C>0$ such that
$\sum_{m=N+1}^{\infty}\frac{m^{m-1}e^{-m}}{m!}<\frac{C}{N^{\frac{1}{2}}},\ N\in\mathbb{N}.$
Since $\{\frac{m^{m-1}e^{-m}}{m!}\}_{m=1}^{\infty}$ is a probability distribution, it then follows that
(6.3) $\sum_{m=1}^{N}\frac{m^{m-1}e^{-m}}{m!}\geq 1-\frac{C}{N^{\frac{1}{2}}},\ N\in\mathbb{N}.$
Using (6.3) for the inequality below, we have
(6.4) $\sum_{\stackrel{{\scriptstyle 1\leq m_{r}\leq N}}{{r=1,\cdots,j_{n}}}}\prod_{r=1}^{N}\frac{m_{r}^{m_{r}-1}e^{-m_{r}}}{m_{r}!}=\left(\sum_{m=1}^{N}\frac{m^{m-1}e^{-m}}{m!}\right)^{j_{n}}\geq\left(1-\frac{C}{N^{\frac{1}{2}}}\right)^{j_{n}}.$
Substituting $N_{n}$ for $N$ in (6.4), a standard estimate gives
(6.5) $\lim_{n\to\infty}\sum_{\stackrel{{\scriptstyle 1\leq m_{r}\leq N_{n}}}{{r=1,\cdots,j_{n}}}}\prod_{r=1}^{N_{n}}\frac{m_{r}^{m_{r}-1}e^{-m_{r}}}{m_{r}!}=1,\ \text{if}\ j_{n}=o(N_{n}^{\frac{1}{2}}).$
We can express $P_{n}^{\text{park}}(\sigma^{-1}_{n+1-j_{n}}<\sigma^{-1}_{n+1-j_{n}+1}<\cdots<\sigma^{-1}_{n})$ as
$P_{n}^{\text{park}}(\sigma^{-1}_{n+1-j_{n}}<\sigma^{-1}_{n+1-j_{n}+1}<\cdots<\sigma^{-1}_{n})=$
$\sum_{\stackrel{{\scriptstyle m_{r}\geq 1;r=1,\cdots,j_{n}}}{{\sum_{r=1}^{j_{n}}m_{r}\leq n}}}P_{n}^{\text{park}}\left(n+1-\sigma^{-1}_{n-j_{n}+l}=\sum_{r=l}^{j_{n}}m_{r},\ l=1,\cdots j_{n}\right).$
Using this with (6.1), we have
(6.6) $P_{n}^{\text{park}}(\sigma^{-1}_{n+1-j_{n}}<\sigma^{-1}_{n+1-j_{n}+1}<\cdots<\sigma^{-1}_{n})=$
$\sum_{\stackrel{{\scriptstyle m_{r}\geq 1;r=1,\cdots,j_{n}}}{{\sum_{r=1}^{j_{n}}m_{r}\leq n}}}\left(\prod_{r=1}^{j_{n}}\frac{m_{r}^{m_{r}-1}e^{-m_{r}}}{m_{r}!}\right)A(n,j_{n};m_{1}\cdots,m_{j_{n}})B(n,j_{n};m_{1},\cdots,m_{j_{n}}).$
If $m_{r}\leq N_{n}$ for $r=1,\cdots,j_{n}$, then trivially, $\sum_{r=1}^{j_{n}}m_{r}\leq j_{n}N_{n}$. Thus, if we choose $j_{n}$ and $N_{n}$ such that
(6.7) $\lim_{n\to\infty}A(n,j_{n};m_{1}\cdots,m_{j_{n}})=1,\ \text{uniformly over}\ \{1\leq m_{r}\leq N_{n};r=1,\cdots,j_{n}\};$
$\lim_{n\to\infty}B(n,j_{n};m_{1},\cdots,m_{j_{n}})=1,\ \text{uniformly over}\ \{1\leq m_{r}\leq N_{n};r=1,\cdots,j_{n}\},$
and such that
(6.8) $j_{n}N_{n}\leq n;\ \text{and}\ j_{n}=o(N_{n}^{\frac{1}{2}}),$
then it will follow from (6.5)-(6.7) that
(6.9) $\lim_{n\to\infty}P_{n}^{\text{park}}(\sigma^{-1}_{n+1-j_{n}}<\sigma^{-1}_{n+1-j_{n}+1}<\cdots<\sigma^{-1}_{n})=1.$
In the calculations below, we always assume that $1\leq m_{r}\leq N_{n}$, for $r=1,\cdots,j_{n}$. We first consider $A(n,j_{n};m_{1}\cdots,m_{j_{n}})$ in (6.2). From Taylor’s remainder formula, we have
(6.10) $-x\geq\log(1-x)\geq-x-\frac{1}{2}x^{2}\frac{1}{(1-x)^{2}},\ x\in(0,1).$
Using (6.10), we have
(6.11) $-\frac{n-1}{n+1}\sum_{r=1}^{j_{n}}m_{r}\geq\log\left(1-\frac{\sum_{r=1}^{j_{n}}m_{r}}{n+1}\right)^{n-1}\geq$
$-\frac{n-1}{n+1}\sum_{r=1}^{j_{n}}m_{r}-(n-1)\left(\frac{1}{2}\frac{(\sum_{r=1}^{j_{n}}m_{r})^{2}}{(n+1-\sum_{r=1}^{j_{n}}m_{r})^{2}}\right).$
From (6.2) and (6.11), we have
(6.12) $\frac{2}{n+1}\sum_{r=1}^{j_{n}}m_{r}\geq\log A(n,j_{n};m_{1}\cdots,m_{j_{n}})\geq$
$\frac{2}{n+1}\sum_{r=1}^{j_{n}}m_{r}-\frac{(n-1)(\sum_{r=1}^{j_{n}}m_{r})^{2}}{2(n+1-\sum_{r=1}^{j_{n}}m_{r})^{2}}.$
Recalling that $\sum_{r=1}^{j_{n}}m_{r}\leq j_{n}N_{n}$, it follows from (6.12) that
(6.13) (6.7)$\ \text{holds for}\ A(n,j_{n};m_{1}\cdots,m_{j_{n}})\ \text{if}\ j_{n}N_{n}=o(n^{\frac{1}{2}}).$ ) .
We now turn to $B(n,j_{n};m_{1}\cdots,m_{j_{n}})$ in (6.2). We write
(6.14) $B(n,j_{n};m_{1}\cdots,m_{j_{n}})=C(n,j_{n};m_{1}\cdots,m_{j_{n}})D(n,j_{n};m_{1}\cdots,m_{j_{n}}),$
where
(6.15) $C(n,j_{n};m_{1}\cdots,m_{j_{n}})=\frac{(n-j_{n})!}{(n-\sum_{r=1}^{j_{n}}m_{r})!}n^{j_{n}-\sum_{r=1}^{j_{n}}m_{r}};$
$D(n,j_{n};m_{1}\cdots,m_{j_{n}})=\frac{\prod_{r=1}^{j_{n}}(n+1-\sum_{r=l}^{j_{n}}m_{r})}{(n+1-\sum_{r=1}^{j_{n}}m_{r})^{\sum_{r=1}^{j_{n}}m_{r}}}n^{\sum_{r=1}^{j_{n}}m_{r}-j_{n}}.$
Let $w_{n}=\sum_{r=1}^{j_{n}}m_{r}$. We suppress the dependence of $w_{n}$ on $\{m_{r}\}_{r=1}^{j_{n}}$. For any choice of $\{m_{r}\}_{r=1}^{j_{n}}$, we have $j_{n}\leq w_{n}\leq j_{n}N_{n}$.
We have
$(n-w_{n})^{w_{n}-j_{n}}n^{j_{n}-w_{n}}\leq C(n,j_{n};m_{1}\cdots,m_{j_{n}})\leq(n-j_{n})^{w_{n}-j_{n}}n^{j_{n}-w_{n}},$
or equivalently,
(6.16) $\left(1-\frac{w_{n}}{n}\right)^{w_{n}-j_{n}}\leq C(n,j_{n};m_{1}\cdots,m_{j_{n}})\leq\left(1-\frac{j_{n}}{n}\right)^{w_{n}-j_{n}}.$
Since $w_{n}\leq j_{n}N_{n}$, it follows from (6.16) that
(6.17) $\text{if}\ j_{n}N_{n}=o(n^{\frac{1}{2}}),\ \text{then}$
$\lim_{n\to\infty}C(n,j_{n};m_{1}\cdots,m_{j_{n}})=1,\ \text{uniformly over}\ \{1\leq m_{r}\leq N_{n};r=1,\cdots,j_{n}\}.$
We have
$\frac{(n+1-w_{n})^{j_{n}}}{(n+1-w_{n})^{w_{n}}}n^{w_{n}-j_{n}}\leq D(n,j_{n};m_{1}\cdots,m_{j_{n}})\leq\frac{n^{j_{n}}}{(n+1-w_{n})^{w_{n}}}n^{w_{n}-j_{n}},$
or equivalently,
(6.18) $\left(1-\frac{w_{n}-1}{n}\right)^{j_{n}}\left(1+\frac{w_{n}-1}{n+1-w_{n}}\right)^{w_{n}}\leq D_{n}\leq\left(1+\frac{w_{n}-1}{n+1-w_{n}}\right)^{w_{n}}.$
Since $w_{n}\leq j_{n}N_{n}$, it follows from (6.18) that
(6.19) $\text{if}\ j_{n}N_{n}=o(n^{\frac{1}{2}}),\ \text{then}$
$\lim_{n\to\infty}D(n,j_{n};m_{1}\cdots,m_{j_{n}})=1,\ \text{uniformly over}\ \{1\leq m_{r}\leq N_{n};r=1,\cdots,j_{n}\}.$
From (6.14), (6.17) and (6.19), it follows that
(6.20) (6.7)$\ \text{holds for}\ B(n,j_{n};m_{1}\cdots,m_{j_{n}})\ \text{if}\ j_{n}N_{n}=o(n^{\frac{1}{2}}).$ ) .
From (6.13), (6.20) and (6.8), we conclude that (6.9) holds if $j_{n}N_{n}=o(n^{\frac{1}{2}})$ and $j_{n}=o(N_{n}^{\frac{1}{2}})$. Thus, if $j_{n}=o(n^{\frac{1}{6}})$ and $N_{n}=[n^{\frac{1}{3}}]$, then (6.9) holds. This completes the proof of the theorem. $\square$
7. Proof of Proposition 1
Let $\sigma=T_{n}(\pi)$, where $T_{n}$ is as in the paragraph containing equation (1.1). Then $\{\sigma_{1},\cdots,\sigma_{m}\}=[m]$ if and only if $\pi_{1}\cdots\pi_{m}\in\mathcal{P}_{m}$. Thus, there are $(m+1)^{m-1}$ choices for $\pi_{1}\cdots\pi_{m}$. Given $\pi_{1}\cdots\pi_{m}\in\mathcal{P}_{m}$, we now consider how many sequences $\pi_{m+1}\cdots\pi_{n}$ there are so that the concatenated sequence $\pi_{1}\cdots\pi_{m}\pi_{m+1}\cdots\pi_{n}$ belongs to $\mathcal{P}_{n}$. Of course, we start with the restriction $1\leq\pi_{j}\leq n$, for all $j\in\{m+1,\cdots,n\}$. It is easy to see that such a sequence $\pi_{m+1}\cdots\pi_{n}$ will be such that the above concatenated sequence belongs to $\mathcal{P}_{n}$ if and only if this sequence results in all $n-m$ cars being able to park in the following scenario: There is a one-way street with $n$ spaces, but with the first $m$ of them already taken up by a trailer. A sequence of $n-m$ cars enters, each with a preferred parking space between 1 and $n$. It is known that the number of such sequences resulting in all $n-m$ cars successfully parking is equal to $(m+1)(n+1)^{n-m-1}$ [2]. Thus, the number of parking functions $\pi\in\mathcal{P}_{n}$ such that $\sigma=T_{n}(\pi)$ satisfies $\{\sigma_{1},\cdots,\sigma_{m}\}=[m]$ is equal to $(m+1)^{m-1}(m+1)(n+1)^{n-m-1}$. Consequently,
$P_{n}^{\text{\rm park}}(\{\sigma_{1},\cdots,\sigma_{m}\}=[m])=\frac{(m+1)^{m-1}(m+1)(n+1)^{n-m-1}}{(n+1)^{n-1}}=\left(\frac{m+1}{n+1}\right)^{m}.$
$\square$